# CaO2: Rectifying Inconsistencies in Diffusion-Based Dataset Distillation

### v1

@@ 1 Introduction
Figure 1: Comparison of validation accuracy and distillation time for different methods under different IPCs on ImageWoof. Our two-stage method is more efficient and can obtain better performance compared with other SOTA approaches. Metrics are averaged across ResNet18, ResNet50, and ResNet101. Detailed results are provided in Appendix 8.
@@ 4.5 Visualization Comparison
We visualize and compare obtained samples using the same random seeds in Fig. 4, qualitatively showing improvements of our method. Our distilled images exhibit higher visual quality and are more discriminative. Interestingly, the changes introduced by the latent optimization process are imperceptible to the human eye. By examining the absolute differences between images before and after latent optimization, we find that the optimization primarily refines object boundaries and background details, suggesting that images might be enhanced by emphasized semantic features and improved robustness against adversarial pixels [4]. Appendix 10 provides more visualizations and analysis.
@@ 4.6 Extending to Different Backbones
Integration with MAR. As discussed in Sec. 3.5, we also extend our method to masked autoregressive generation models [19]. Specifically, we adopt the MAR-Base model, generating images by sampling 32 autoregressive steps with 100 denoising steps per autoregressive step. We select the correct and most confident samples from the image pool of size 2$\times$×IPC. For latent optimization, we adopt a maximum token masking ratio of 0.25, train for 100 iterations with a learning rate of 0.0001, and also use the $L_{\infty}$ norm for regularization. Tab. 5 presents the results showing that MAR alone is already a robust baseline for performance improvement on ImageNet subsets. When combined with our method, it achieves an average accuracy increase of 3.3% on ImageWoof and 2.0% on ImageNette across two IPC settings. Visualization results are provided in Appendix 10.
@@
Integration with Minimax Diffusion. We further integrate our method with Minimax Diffusion [9]. Minimax Diffusion finetunes the diffusion transformer model to enhance the representativeness and diversity of generated samples without changing the model architecture. In our setup, we select from an image pool of size 2$\times$×IPC, apply the $L_{\infty}$ norm, and use the true label as guidance, except for ImageWoof at IPC=10, where we use zero label embeddings instead. Tab. 5 shows that we can further improve the accuracy upon the Minimax Diffusion method, with an average improvement of 4.3% on ImageWoof and 2.0% on ImageNette. Visualizations are available in Appendix 10.
In summary, the results demonstrate that CaO2 functions effectively as a plug-and-play module for diffusion models with diverse training objectives, as well as for masked autoregressive models that use a distinct generation scheme compared to conventional diffusion models. More ablations and discussions are included in Appendix 8.

### v2

@@ 1 Introduction
Figure 1: Comparison of validation accuracy and distillation time for different methods under different IPCs on ImageWoof. Our two-stage method is more efficient and can obtain better performance compared with other SOTA approaches.
@@ 4.5 Visualization Comparison
We visualize obtained samples using the same random seeds in Fig. 4, qualitatively revealing that our method produces more discriminative and visually refined images. Interestingly, latent optimization introduces imperceptible changes to the human eye. By examining the absolute differences between images before and after latent optimization, we find that the optimization primarily refines object boundaries and background details, suggesting that images might be enhanced by emphasized semantic features and improved robustness against adversarial pixels [4]. Appendix 10 provides more visualizations and analysis.
@@ 4.6 Extending to Different Backbones
Integration with MAR. As discussed in Sec. 3.5, we extend our method to masked autoregressive models [19], using MAR-Base with 32 autoregressive steps and 100 denoising steps per autoregressive step. From a 2$\times$×IPC image pool, we select the most confident correct samples. For latent optimization, we apply a 0.25 max masking ratio, train for 100 steps with a 0.0001 learning rate, and use $L_{\infty}$ regularization. Tab. 5 shows that MAR alone is a strong baseline on ImageNet subsets, and combining it with our method boosts accuracy by 3.3% on ImageWoof and 2.0% on ImageNette. Visualizations are in Appendix 10.
@@
Integration with Minimax Diffusion. We further integrate our method with Minimax Diffusion [9], a finetuned diffusion transformer that improves sample quality without altering the model architecture. Using a 2$\times$×IPC image pool, $L_{\infty}$ regularization, and true label guidance (except zero embeddings for ImageWoof at IPC=10), Tab. 5 shows that we further improve the performance with an average improvement of 4.3% on ImageWoof and 2.0% on ImageNette. Visualizations are available in Appendix 10. In summary, CaO2 serves as an effective plug-and-play module for both diffusion models with diverse training objectives and masked autoregressive models with distinct generation schemes.
@@ 5 Conclusion
Acknowledgments: This research is supported by NSF IIS-2525840, CNS-2432534, ECCS-2514574, NIH 1RF1MH133764-01 and Cisco Research unrestricted gift. This article solely reflects opinions and conclusions of authors and not funding agencies.